# SynSUM – Synthetic Benchmark with Structured and Unstructured Medical Records

### v1

@@ Abstract
We present the SynSUM benchmark, a synthetic dataset linking unstructured clinical notes to structured background variables. The dataset consists of 10,000 artificial patient records containing tabular variables (like symptoms, diagnoses and underlying conditions) and related notes describing the fictional patient encounter in the domain of respiratory diseases. The tabular portion of the data is generated through a Bayesian network, where both the causal structure between the variables and the conditional probabilities are proposed by an expert based on domain knowledge. We then prompt a large language model (GPT-4o) to generate a clinical note related to this patient encounter, describing the patient symptoms and additional context. The SynSUM dataset is primarily designed to facilitate research on clinical information extraction in the presence of tabular background variables, which can be linked through domain knowledge to concepts of interest to be extracted from the text - the symptoms, in the case of SynSUM. Secondary uses include research on the automation of clinical reasoning over both tabular data and text, causal effect estimation in the presence of tabular and/or textual confounders, and multi-modal synthetic data generation. The dataset can be downloaded from https://github.com/prabaey/SynSUM.
@@ 1 Introduction
Electronic health records (EHRs) are a gold mine of information, containing a mix of structured tabular variables (medication, diagnosis codes, lab results…) and free unstructured text (detailed clinical notes from physicians, nurses…) (Ford et al., 2016). These EHRs form a valuable basis for training clinical decision support systems, (partially) automating essential processes in the clinical world, such as diagnosis, writing treatment plans, and more (Peiffer-Smadja et al., 2020; Mujtaba et al., 2019; Rasmy et al., 2021; Li et al., 2020; Xu et al., 2019). While large language models can help leverage the potential of the unstructured text portion of the EHR (Zhang et al., 2020; Liu et al., 2022; Huang, Altosaar, and Ranganath, 2019; Lehman and Johnson, 2023; Singhal et al., 2023; Labrak et al., 2024), these black box systems lack interpretability (Quinn et al., 2022; Zhao et al., 2024; Tian et al., 2024). In high-risk clinical applications, it can be argued that one should prefer more robust and transparent systems built on simpler, feature-based models, like regression models, decision trees, Bayesian networks, etc (Rudin, 2019; Sanchez et al., 2022; Lundberg et al., 2020). However, such models cannot directly deal with unstructured text and require tabular features as an input. For this reason, automated clinical information extraction (CIE) (Ford et al., 2016; Wang et al., 2018; Hahn and Oleynik, 2020) remains an essential tool for building large structured datasets that can serve as training data for such systems.
However, CIE remains a challenging task due to the complex nature of clinical notes. These often leave out important contextual details that an inexperienced reader would not know about, and which an automated system would need in order to correctly extract concepts from the text. Existing systems do not fully exploit the available medical domain knowledge to fill in this gap. We propose that clinical information extraction could benefit from leveraging two additional sources of information, apart from the unstructured text itself. On the one hand, we have the tabular features already encoded in the EHR, containing information on the medical history of the patient, as well as partially encoded information relating to the current visit. On the other hand, we can connect this encoded background information with the concepts we are trying to extract from the text, using medical domain knowledge. Such domain knowledge could be structured in the form of a Bayesian network. This idea is presented conceptually in Figure 1.
To put this idea into practice, we need a clinical dataset which (i) contains a mix of tabular data and unstructured text, where (ii) the tabular concepts and the concepts we aim to extract from the text can be linked through domain knowledge. While open-source datasets like MIMIC-III and MIMIC-IV contain this mix, they are too complex for the preliminary research we envision. First, the area of intensive care in which the data was collected is very extensive, making it hard to isolate a specific small-scale use-case for which the domain knowledge could be listed. Second, the portion of the dataset which is encoded into tabular features is mostly driven by billing needs, rather than completeness or accuracy, and does not contain any encoded symptoms, which are the concepts we would like to extract from the text. Third, the link between the tabular features and the concepts mentioned in the text might be inconsistent due to system design or human errors, as shown by Kwon et al. (2024). Finally, the EHRs in MIMIC are time series, adding another layer of complexity.
In this work, we build a synthetic yet sufficiently realistic dataset that addresses some of these shortcomings and allows research on the idea of incorporating domain knowledge for improved CIE in the presence of encoded tabular variables. Our dataset, called SynSUM (Synthetic Structured and Unstructured Medical records) is a self-contained set of synthetic EHRs in a primary care setting, fulfilling the following requirements:
  • There is a mix of structured tabular data and unstructured text.
  • The use-case in which the EHR is constructed allows us to explicitly model the domain knowledge in the form of a Bayesian network, which links the background tabular variables with the concepts expressed in the text.
  • There is no time aspect, we construct the EHRs to be a static snapshot of a single patient encounter.
As a specific use-case, we choose the domain of respiratory diseases in primary care. Specifically, we envision the scenario where a patient visits a primary care doctor with some respiratory symptoms. The doctor notes down the patient’s symptoms in a clinical note, along with some additional context, and stores this in the electronic patient record together with the encoded diagnosis, as well as the encoded symptoms. In a realistic setting, the primary care doctor would only encode the primary complaint, rather than all symptoms. However, since we aim to build a dataset which can be used to train CIE models, where the concepts of interest are the symptoms expressed in the text, we encode all these symptoms in the artificial EHR. The tabular portion of the EHR additionally contains background information on the underlying health conditions of the patient, which may help the doctor with their diagnosis.
Section 2 describes the process of generating our SynSUM dataset of synthetic patient records in the domain of respiratory diseases. This dataset is made up of 10.000 fictional patient encounters, each represented by a set of structured tabular variables and free clinical text notes describing the encounter. Then, Section 3 runs some simple symptom predictors on our dataset, which can form a baseline for further research. Finally, Section 4 discusses the potential application of our dataset to clinical information extraction, as well as to other areas of research, while also highlighting its limitations.
2 Data generation
Figure 2: Overview of the full data generating process for the SynSUM dataset. First, the tabular portion of the synthetic patient record is sampled from a Bayesian network, where both the structure and the conditional probability distributions were defined by an expert. Afterwards, we construct a prompt containing information on the symptoms experienced by the patient, as well as their underlying health conditions (but no diagnoses). We ask the GPT-4o large language model to generate a fictional clinical note describing this patient encounter. Finally, we ask to generate a more challenging compact version of the note, mimicking the complexity of real clinical notes by prompting the use of abbreviations and shortcuts. We generate 10.000 of these synthetic patient records in total.

### v2

@@ Abstract
We present the SynSUM benchmark, a synthetic dataset linking unstructured clinical notes to structured background variables. The dataset consists of $10{,}000$ artificial patient records containing tabular variables (like symptoms, diagnoses and underlying conditions) and related notes describing the fictional patient encounter in the domain of respiratory diseases. The tabular portion of the data is generated through a Bayesian network, where both the causal structure between the variables and the conditional probabilities are proposed by an expert based on domain knowledge. We then prompt a large language model (GPT-4o) to generate a clinical note related to this patient encounter, describing the patient symptoms and additional context. We conduct both an expert evaluation study to assess the quality of the generated notes, as well as running some simple predictor models on both the tabular and text portions of the dataset, forming a baseline for further research. The SynSUM dataset is primarily designed to facilitate research on clinical information extraction in the presence of tabular background variables, which can be linked through domain knowledge to concepts of interest to be extracted from the text - the symptoms, in the case of SynSUM. Secondary uses include research on the automation of clinical reasoning over both tabular data and text, causal effect estimation in the presence of tabular and/or textual confounders, and multi-modal synthetic data generation.
Code & Datasethttps://github.com/prabaey/SynSUM
@@ 1 Introduction
Electronic health records (EHRs) are a gold mine of information, containing a mix of structured tabular variables (medication, diagnosis codes, lab results…) and free unstructured text (detailed clinical notes from physicians, nurses…) (Ford et al. 2016). These EHRs form a valuable basis for training clinical decision support systems, (partially) automating essential processes in the clinical world, such as diagnosis, writing treatment plans, and more (Peiffer-Smadja et al. 2020; Mujtaba et al. 2019; Rasmy et al. 2021; Li et al. 2020; Xu et al. 2019). While large language models can help leverage the potential of the unstructured text portion of the EHR (Zhang et al. 2020; Liu et al. 2022; Huang, Altosaar, and Ranganath 2019; Lehman and Johnson 2023; Singhal et al. 2023; Labrak et al. 2024), these black box systems lack interpretability (Quinn et al. 2022; Zhao et al. 2024; Tian et al. 2024). In high-risk clinical applications, it can be argued that one should prefer more robust and transparent systems built on simpler, feature-based models, like regression models, decision trees, or Bayesian networks (Rudin 2019; Sanchez et al. 2022; Lundberg et al. 2020). However, such models cannot directly deal with unstructured text and require tabular features as an input. For this reason, automated clinical information extraction (CIE) (Ford et al. 2016; Wang et al. 2018; Hahn and Oleynik 2020) is an essential tool for building large structured datasets that can serve as training data for such systems.
However, CIE remains a challenging task due to the complex nature of clinical notes. These often leave out important contextual details which an automated system would need in order to correctly extract concepts from the text. Existing systems do not fully exploit the available medical domain knowledge to fill in this gap. We propose that CIE could benefit from leveraging two additional sources of information, apart from the unstructured text itself. On the one hand, a range of tabular features are already encoded in the EHR. These contain information related to a particular patient visit (e.g. partially encoded symptoms or diagnosis codes), as well as information on the medical history of the patient. On the other hand, we can connect this encoded background information with the concepts we are trying to extract from the text, using a Bayesian network (BN) that represents medical domain knowledge. A visual example that helps illustrate this idea is shown in Figure 2 in Appendix A.
To investigate and implement the described research idea, we need a clinical dataset which (i) contains a mix of tabular data and unstructured text, where (ii) the tabular data and the concepts we aim to extract from the text can be linked through domain knowledge. While open-source datasets like MIMIC-III (Johnson et al. 2016) and MIMIC-IV (Johnson et al. 2023) contain this mix, they are not a perfect fit. First, the area of intensive care in which the data was collected is very extensive, making it hard to isolate a specific small-scale use-case for which the domain knowledge could be listed. Second, the portion of the dataset which is encoded into tabular features is often driven by billing needs, rather than completeness or accuracy, and does not contain any encoded symptoms, which are concepts that could be interesting to extract from the text for application in clinical decision support systems. Third, the link between the tabular features and the concepts mentioned in the text might be inconsistent due to system design or human errors (Kwon et al. 2024). Finally, the EHRs in MIMIC being time series adds another layer of complexity. Other existing datasets linking unstructured clinical text to structured features include BioDEX (D’Oosterlinck et al. 2023), a large set of papers describing adverse drug events, as well as TCGA-Reports (Kefeli and Tatonetti 2024), a set of cancer pathology reports, both accompanied by tabular patient descriptors and extracted biological features. However, in both cases, it is not trivial to devise a BN representing the relevant expert knowledge, partially because that knowledge is not fully understood yet.
In this work, we build a synthetic yet sufficiently realistic dataset that addresses some of these shortcomings, enabling research on incorporating domain knowledge for improved CIE in the presence of tabular variables. Our dataset, called SynSUM (Synthetic Structured and Unstructured Medical records) is a self-contained set of synthetic EHRs in a primary care setting, fulfilling the following requirements:
  • It mixes structured tabular data and unstructured text.
  • By design, clinical concepts expressed in the text and encoded in the tabular portion of the dataset are connected through a Bayesian network representing domain knowledge. In this case, the domain is respiratory diseases, with their associated symptoms and underlying conditions.
  • Each EHR is a static snapshot of a single patient encounter, eliminating the time aspect for simplicity.
SynSUM is constructed in the domain of respiratory diseases, simulating patient visits to a primary care doctor. We mimic the scenario where the doctor notes down the patient’s symptoms in a clinical note, along with some additional context, and stores this in the EHR together with the encoded diagnosis, as well as the encoded symptoms. Additionally, the EHR stores tabular background information on the underlying health conditions of the patient.
Section 2 describes how we generated 10,000 fictional patient encounters by sampling from an expert-defined BN and prompting a large language model. Section 3 describes the expert evaluation conducted by a panel of five primary care physicians on a subset of the generated notes. Section 4 then runs some simple symptom predictors on both the tabular and textual portions of our dataset, which form a baseline to be exceeded by future research. Finally, Section 5 discusses the potential use of our dataset in various areas of research (including but not limited to CIE), while also highlighting its limitations.
2 Data generation
Figure 1: Overview of the full data generating process for the SynSUM dataset. First, the tabular portion of the synthetic patient record is sampled from a Bayesian network, where both the structure and the conditional probability distributions were defined by an expert. Afterwards, we construct a prompt describing the symptoms experienced by the patient, as well as their underlying health conditions (but no diagnoses). We ask the GPT-4o large language model to generate a clinical note describing this patient encounter. Finally, we ask to generate a more challenging compact version of the note, mimicking the complexity of real clinical notes by prompting the use of abbreviations and shortcuts. We generate $10{,}000$ of these synthetic patient records in total.
Our general methodology for generating the artificial patient records is shown in Figure 1. We now zoom in on the two major parts of this data generating process. First, Section 2.1 describes how we generated the structured tabular variables through an expert-defined Bayesian network (BN). Then, Section 2.2 dives into the clinical note generation by the large language model.
2.1 Modeling structured tabular variables with a Bayesian network
Causal structure
We asked an expert to define a Directed Acyclic Graph (DAG) which (partially) models the domain of respiratory diseases in primary care, shown in Figure 1. In this DAG, a directed arrow between two variables models a causal relation between them. Central to the model are the diagnoses of pneumonia and common cold, which may give rise to five symptoms (dyspnea, cough, pain, fever and nasal). The expert also modeled some relevant underlying conditions which may render a patient more predisposed to certain diagnoses or symptoms: asthma, smoking, COPD and hay fever. Based on the symptoms experienced by a patient, a primary care doctor decides whether to prescribe antibiotics or not. The presence and severity of the symptoms, as well as the prescription of antibiotics as a treatment, influence the outcome, which is the total number of days (# days) that the patient eventually stays home as a result of illness. Finally, there are some non-clinical variables which exert an external influence on the diagnoses, treatment and outcome (season, policy and self-employed). Table 3 in Appendix B summarizes all variables and their meaning, as well as their possible values. While this model does not completely describe the real world, we do believe it to be sufficiently realistic for the purpose of generating a useful artificial dataset.
Probability distribution
We turn the DAG into a BN by defining a joint distribution, which factorizes into $16$ conditional distributions, one for each variable. We use four different approaches to define these conditionals, after which we can sample synthetic patients from the BN in a top-down fashion.
For each of the variables asthma, smoking, hay fever, COPD, season, pneumonia, common cold, fever, policy and self-employed, we ask our expert to define a conditional probability table (CPT). When a variable has many parents, it becomes infeasible to manually fill in the CPT in a clinically meaningful way. For this reason, we use a Noisy-OR distribution for the symptoms dyspnea, cough, pain and nasal, which assumes an independent causal mechanism behind the activation of a symptom through any of its parents. We model the prescription of antibiotics through a logistic regression model, mimicking the way the clinician’s suspicion of pneumonia (which needs to be treated with antibiotics) rises when a higher number of symptoms is present in the patient, with some symptoms weighing more than others, eventually exceeding a threshold. These weights are defined by the expert and validated with a set of test cases. In a similar fashion, we model the # days the patient stays home as a result of their complaints with a Poisson regression model. These weights are learned through maximum likelihood estimation over a set of train cases provided by the expert. The full specification of our BN can be found in Appendix B.
2.2 Generating unstructured text with a large language model
Once the tabular patient record has been generated, we prompt a large language model (LLM, in our case GPT-4o) to write a clinical note based on the tabular variables associated with this fictional encounter. For this, we use a prompt which is generally structured like the example shown in Figure 1 (except for some special cases, which are addressed in Appendix C). Additional example prompts can be found in Appendix D. Only the background variables and symptoms may be directly mentioned in the prompt, while the diagnoses pneumonia and common cold would not yet be known to a clinician who is taking notes during a consultation and are therefore left out. They can still influence the content of the note through the descriptions of the symptoms that are included in the prompt. The treatment and outcome are left out of the prompt as well, just like the non-clinical variables, as all of these are typically either unknown or irrelevant at the time of writing the note.We ask the LLM to generate both a “history” portion and a “physical examination” portion, to encourage variation and complexity of the generated notes.
In general, the prompt is constructed as follows. For more details, we refer to Appendix C. First, we list the symptoms which are experienced by the patient. We do not exhaustively list the full set of symptoms, but rather sample the probability that a symptom is mentioned according to a distribution defined by the expert. This renders the notes more realistic, since real notes do not exhaustively mention the presence or absence of all possible symptoms, but rather follow the narrative of the patient and the subsequent probing of the clinician. While we encourage the LLM to invent additional context around the patient’s symptoms, we want this to at least partially relate to the underlying cause of the symptom. For this reason, we include a descriptor next to each symptom, which is randomly sampled from a set of expert-defined phrases that describe the symptom in the case where it results from a particular cause. This way, the diagnoses can influence the content of the note, even when they are not mentioned explicitly in the prompt. The full list of descriptors, as well as further information on how we sample them, is included in Appendix C. The next part of the prompt lists the underlying health conditions, which are assumed to be known up-front, and could therefore influence the content of the note. Finally, the prompt lists some additional instructions, which we motivate in Appendix C. There, we also outline an alternative prompting strategy, which is used to generate clinical notes for the patient records where none of the respiratory symptoms are present, which is the case for around one third of our dataset. In these cases, we encourage the LLM to imagine an alternative reason for the patient visit, unrelated to the respiratory domain.
Real clinical notes can be challenging, often containing abbreviations, shortcuts and denser sentence structure. To make the notes more challenging, we ask the LLM to create a compact version of each note through an additional prompt, as can be seen in Figure 1. Our dataset contains both the original note and the compact version of the note.
3 Expert evaluation
Normal Compact
Consis- tency Realism (hist) Realism (phys) Clinical accuracy Content Reada- bility
mean 4.69 4.53 4.15 4.92 4.88 4.02
std 0.12 0.21 0.30 0.07 0.10 0.31
Table 1: Results of the expert evaluation study. We report average scores (ranging from 1-5) over $30$ notes, together with their standard deviation over the five evaluators.
We asked five general practitioners to evaluate the quality of a random sample of $30$ generated notes. The evaluators got to see the two versions of the generated note (normal and compact), as well as the prompt that was used to generate them. They were asked to evaluate four aspects:
  1. 1.
    Consistency: The description of the patient’s symptoms and underlying health conditions must correspond with the instructions provided in the prompt.
  2. 2.
    Realism: The context and details invented by the LLM and added to the “history” portion of the note should be realistic given the symptoms experienced by the patient, as detailed in the prompt. Furthermore, the elements that are checked in the “physical examination” must be realistic in light of the patient information described in the “history”. We evaluate both aspects separately, focusing on content (what information is included in the note), rather than format (how the information is written down).
  3. 3.
    Clinical accuracy: The findings described in the “physical examination” must be clinically accurate, both in a standalone fashion and in relation to the patient’s symptoms described in the “history”.
  4. 4.
    Quality of compact version: The content of the compact version of the note must correspond well with the original note, and remain readable despite the use of abbreviations and shortcuts. We evaluate both aspects separately.
To measure consistency, the evaluators were asked to assign a penalty for every element in the note that does not correspond with the requested information in the prompt, and for every requested symptom that was missing from the generated note. These penalties were then turned into ratings from 1 (>3 penalties) to 5 (0 penalties). All other aspects were directly rated on a scale of 1 (very bad) to 5 (perfect). Table 1 shows the results. For each rated aspect, we calculate the average score over all $30$ notes, and report the mean and standard deviation over the five evaluators. For more details on the meaning of each rating in each aspect, as well as inter-annotator agreement scores, we refer to Appendix E.
Our artificial notes were rated as highly consistent with the prompt, and therefore with the information present in the tabular portion of the dataset. As shown in Appendix E.2, a large majority of the inconsistencies arise from a violation of the additional instructions (usually by inventing additional symptoms), while the key information included in the prompt was still conveyed correctly in the note.
The evaluators also deem the notes sufficiently realistic, though the realism of the history section is rated higher on average than the realism of the physical examination. Out of those notes that scored worse, many included a clinical test that seemed unnecessary to the evaluators, while a few forgot a test that was deemed important. At the same time, the very high score for clinical accuracy is an important indication that the notes do not contain falsehoods.
Multiple evaluators mentioned that while the content of the notes seemed realistic, the format did not, as their own notes would be more complex as opposed to the artificial notes, which use clean language and full sentences. This underscores the fact that the dataset should not be used to train any systems which will later be deployed on real notes, and should instead fulfill the role of a research benchmark only.
The compact versions of the notes score very well on content, mostly conveying the same information as the original. They score a little lower in terms of readability, which evaluators often attributed to the extensive use of abbreviations.
4 Symptom predictor baselines
In order to set a baseline for future information extraction tasks, we run various prediction models on both the tabular and textual parts of the dataset. These models are trained to predict each of the five symptoms: dyspnea, cough, pain, fever and nasal.
  • $\mathcal{P}(\text{{sympt}}\mid\text{{all}})$: Predict the symptom given all other tabular features as evidence. This set includes the background, diagnoses, non-clinical, treatment and outcome variables, as well as the other symptoms.
[…]
  • $\mathcal{P}(\text{{sympt}}\mid\text{{realistic}})$: Predict the symptom given a more realistic set of tabular features as evidence. We do not expect policy, self-employed and #days to be recorded in any kind of realistic patient record, and therefore leave them out of this evidence set. As in the no-sympt setting, we do not include the symptoms either. In other words, this set includes the background, diagnoses, season and treatment variables.
Apart from the tabular-only baselines, we also train some baselines that get to see the text. Our neural-text classifier takes only the text as an input (in the form of a pretrained clinical sentence embedding) and outputs the probability that a symptom is mentioned in the text. We extend this text-only baseline by concatenating a numerical representation of the tabular features to the text embedding at the input, forming the neural-text-tab baseline. Again, we do this for each of the three evidence settings outlined above. Note that this is the only model that combines both the background knowledge available in the tabular features with the unstructured text, and it does so in a naive way. Future work will focus on improving the performance of this model by exploiting the relations between any of the tabular concepts, as envisioned in Figure 2.
dyspnea cough pain nasal fever
BN-tab
- all 0.7370 0.7816 0.2386 0.7146 0.4864
- no-sympt 0.7153 0.7776 0.1312 0.7146 0.4384
- realistic 0.6698 0.7763 0.0280 0.7146 0.3594
XGBoost-tab
- all 0.6639 0.7848 0.4070 0.7130 0.4111
- no-sympt 0.6612 0.7779 0.3638 0.7146 0.4015
- realistic 0.6626 0.7798 0.3698 0.7146 0.3951
neural-text
- normal 0.9660 0.9595 0.8415 0.9602 0.9125
neural-text-tab
- normal + all 0.9526 0.9481 0.8096 0.9598 0.8804
- normal + no-sympt 0.9592 0.9530 0.8078 0.9550 0.9014
- normal + realistic 0.9544 0.9543 0.8303 0.9575 0.9101
neural-text
- compact 0.9383 0.9480 0.7828 0.9583 0.9073
neural-text-tab
- compact + all 0.9535 0.9384 0.7675 0.9566 0.8987
- compact + no-sympt 0.9363 0.9240 0.7984 0.9638 0.8991
- compact + realistic 0.9442 0.9422 0.7880 0.9606 0.9051
Table 2: F1-score obtained over the test set for each of our baseline models. The results for the text classifiers trained over the normal vs. the compact version of the notes are grouped together for readability. We report results for the mean embedding type, while results for the other embedding types can be found in Table 8. We underline the best result obtained by the tabular-only models, while the best overall result per symptom is in shown in bold.
4.1 Models
BN-tab
We provide the causal structure in Figure 1 to the BN, and learn all parameters in the conditional probability tables (CPTs), Noisy-OR distributions, logistic regression model and Poisson regression model from the training data. In each case, we use maximum likelihood estimation to estimate the parameters. Where we don’t directly learn a CPT (for the variables dyspnea, cough, pain, nasal, antibiotics and #days), we evaluate the learned distribution for each combination of child and parent values to obtain a CPT. For more details, we refer to Appendix F.1. We then use variable elimination over the full joint distribution to evaluate the capability of the learned BN to predict each of the symptoms, taking different variables as evidence according to the three settings described earlier (all, no-sympt and realistic).
XGBoost-tab
We train an XGBoost classifier for each symptom in combination with each of the three evidence settings, meaning each classifier sees a different set of tabular features at the input. We optimize the hyperparameters separately for each combination (15 in total) using 5-fold cross-validation. For more details, we refer to Appendix F.2.
Neural-text
We train a neural classifier that takes only the text as an input and is trained to predict the probability a symptom is mentioned. We train separate classifiers for each symptom. We first split the text into sentences, and transform these into an embedding using the pretrained clinical representation model BioLORD-2023 (Remy, Demuynck, and Demeester 2024). We explore 4 settings for turning these sentence embeddings into a single note embedding:
The note embedding is then fed into a multi-layer perceptron with one hidden layer, followed by a Sigmoid activation for the symptoms dyspnea, cough, pain and nasal, and a Softmax activation with 3 outputs heads for the symptom $fever$. We optimized the parameters of each model (i.e. each combination of symptom and embedding type) using the binary or multiclass cross-entropy objective over the symptom labels. For more details, we refer to Appendix F.3.
Neural-text-tab
We extend the neural-text baseline by concatenating the mean text embeddings with the tabular variables at the input of each neural classifier. All categorical tabular variables were first transformed to a one-hot encoding, while the variable #days was preprocessed using standard scaling. We used the same architecture as the neural-text baseline (only changing the dimension of the input layer), and again trained separate classifiers for each symptom combined with each evidence setting (all, no-sympt and realistic). For more details, we refer to Appendix F.4.
4.2 Results
We use a random $8{,}000$/$2{,}000$ split to obtain a train and test set. We use cross-validation on the train set to tune any hyperparameters, and report the final F1-score over the test set after training. For the binary symptoms, we use a $0.5$ decision threshold. For fever, which has three possible categories, the class with the highest predicted probability is chosen. In that case, we report the macro F1-score.
Table 2 compares the results obtained for all baselines. The tabular-only baselines (BN-tab and XGBoost-tab) perform consistently worse than the baselines that include text (neural-text and neural-text-tab). The evidence setting where all other features are included as evidence usually performs best for the tabular-only baselines.
The neural-text-tab baseline does not perform better than the neural-text baseline for the normal notes. While there is little room for improvement in the dyspnea, cough and nasal classifiers, the symptoms pain and fever are harder to predict. We also note a consistent gap in performance between the normal and compact notes, which can be attributed to the higher complexity of the latter. In that case, the neural-text-tab classifier manages to marginally improve over the neural-text classifier by including the tabular features.
Table 8 in Appendix F.3 further breaks down the results for the neural-text classifier over the different embedding types. Using only the hist embedding outperforms the phys-only setting for the symptoms cough, pain and fever. While the score for hist comes close to those for mean and concat, the latter usually still outperform the former for the normal notes, showing that there is some complementary information in the “history” and “physical examination” portions.
5 Discussion
Symptom extraction
Our analysis of simple tabular and textual baseline models revealed that the symptoms pain and fever are hardest to predict in both the tabular-only and the text-only setting. Combining both settings, i.e. integrating tabular background features in the extraction of the concepts from the text, and linking them through domain knowledge, may have potential for improved information extraction. Future work will focus on realizing this hybrid approach to improve upon the baseline results presented in Table 2.
Potential uses
The dataset is primarily designed to facilitate research on clinical information extraction in the presence of tabular background variables. Future work will focus on realizing the idea presented in Figure 2, where the tabular features aid in more accurately extracting concepts from the text by linking them through domain knowledge. Apart from this, we also foresee multiple secondary uses of the dataset. First, the dataset could facilitate research on the automation of clinical reasoning over tabular data and text, following the example of Rabaey et al. (2024). Second, it could be used to benchmark causal effect estimation methods in the presence of textual confounders, similar to Veitch, Sridhar, and Blei (2020), thanks to the purposeful inclusion of both a treatment and outcome variable in our dataset. Third, there has been increasing interest in clinical synthetic data (Hernandez et al. 2022), where a set of patient characteristics is turned into a synthetic version that is meant to protect the privacy of individuals in the original dataset. Our dataset could serve as a benchmark for comparing synthetic data generation methods that jointly generate tabular variables and text (Lee 2018; Ceritli et al. 2023; Guan et al. 2019). In short, any area of research focusing on the intersection of tabular data and text in healthcare can potentially benefit from our proposed benchmark.
Limitations
While the dataset we constructed is meant to have realistic properties, we also intentionally simplify reality to make the design and generation process feasible. The dataset is purely meant as a research benchmark where the ground truth relations are known, and results obtained on it are not meant to transfer to real clinical notes or datasets. This is confirmed by the evaluators in our expert study, who warned that the writing style used in the text notes is not in accordance with reality, even if the content is mostly realistic. We therefore advise strongly against using the dataset for training prediction models which will be deployed in real settings. Furthermore, our dataset contains static patients that do not evolve over time. While this is a reasonable simplification to make in the realm of primary care, time series are a crucial modality in more complex settings, such as intensive care.
Acknowledgments
Paloma Rabaey and Henri Arno’s research is funded by the Research Foundation Flanders (FWO Vlaanderen) with grant numbers 1170124N and 11Q2C24N. This research also received funding from the Flemish government under the “Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen” programme. The authors would like to thank Géraldine Deberdt, Thibault Detremerie, An De Sutter, Veerle Piessens en Florian Stul for participating in the expert evaluation of the artificial clinical notes.
Appendix
Appendix A Clinical Information Extraction with Background Knowledge
@@
Clinical information extraction can be improved by exploiting tabular background information that is present in the EHR, next to the text from which we are extracting clinical concepts (e.g. the symptoms). Figure 2 illustrates this idea with two practical examples. The SynSUM benchmark enables future research to explore this idea, by linking structured tabular features with unstructured text describing a fictional patient encounter.
Appendix B Bayesian network